OpenStudy (anonymous):

y'' + 9y = 2xsinx + xe^(3x) Need to solve this using method of undetermined coefficients. What I need to know: coefficients of particular solutions. For example, the "A" coefficient for first term (2xsinx) is 1/4. What is B since in my particular solution for 2xsinx (without the second term -- xe^(3x)) is (Ax+B)cosx+(Ax+B)sinx. Please help!

OpenStudy (anonymous):

\[ \frac{d^2y}{dx^2} + 9 \frac{dy}{dx} = 2xsinx + xe^{3x}\]Just to make it look better. : )

OpenStudy (anonymous):

Ok, thank you. Have any idea for a solution?

OpenStudy (anonymous):

I am trying. : )

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%22+%2B9y%27+%3D+2xsinx+%2Be^%283x%29 Out of my League Sorry.

OpenStudy (anonymous):

I assume you can solve for fundamental solution and only wants particular soloution?

OpenStudy (anonymous):

Yes. I can find the homogenous solution which is : A*cos(3x) + B*sin(3x)

OpenStudy (anonymous):

\[My particular solution: (AX+B)cosx+(AX+B)sinx. After differentiating the particular solution and putting the particular solution along with its derivatives \in the main equation (but with only 2xsinx) I get: 8AXsinx+8Acosx-2Asinx+8Bsinx+8Bcosx+2Acosx=2xsinx. From this I get A=1/4. What is b?\]

OpenStudy (anonymous):

Myparticularsolution:(AX+B)cosx+(AX+B)sinx.After differentiating the particular solution and putting the particular solution along with its derivatives the main equation (but with only 2xsinx) I get: 8AXsinx+8Acosx−2Asinx+8Bsinx+8Bcosx+2Acosx=2xsinx. From this I get A=1/4.What is B? * 2 minutes ago

OpenStudy (anonymous):

You can also just try to solve: \[y \prime \prime + 9y = 2xsinx\] All I need is the coefficient B of the particular solution which is: \[(AX+B)cosx + (AX+B)sinx\] If somebody knows this please help.

myininaya (myininaya):

give me a sec i'm looking at it k?

OpenStudy (anonymous):

ok

myininaya (myininaya):

give me a second and i will check my work

OpenStudy (anonymous):

ok.

myininaya (myininaya):

oops i forgot about that 9

myininaya (myininaya):

you know the one in front of the y i forgot about that one that changes things

OpenStudy (anonymous):

Does the x in xsin x mess things up at all?

OpenStudy (anonymous):

in front of sin x..

myininaya (myininaya):

give me a few more minutes to redo it

OpenStudy (anonymous):

@myininaya Shouldn't you have a x term before cosx, like \[(x/4)cosx\]?

OpenStudy (anonymous):

Maybe (dunno) u have to multiply the trial sol'n by x...

OpenStudy (anonymous):

@estudier. No it doesn't. You make a particular solution to be: (AX+B)cosx + (AX+B)sinx, then differentiate, then put that in the original equation and you should get A and B from it.

OpenStudy (anonymous):

Yes, that might work, I suppose... I'm only used to doing the pieces separately.

OpenStudy (anonymous):

How u doing, myinanaya?

myininaya (myininaya):

okay \[y_p=-\frac{1}{16}\cos(x)+\frac{1}{4}xsin(x)+\frac{-1}{54}e^{3x}+\frac{1}{18}xe^{3x}\]

myininaya (myininaya):

now let me check this lol

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Look: \[8AXsinX+8AcosX-2AsinX+8BsinX+8BcosX+2AcosX=2XsinX\] Now, from this since there is only one term on the left that corresponds to \[XsinX\] \[8A = 2\]\[A = 1/4\] Since there is no terms on the right with \[cosX\] or \[sinX\] we set sum of every coefficient on the left that has corresponding sinX or cosX to 0 (zero). What is B? Is it 0 (zero) or is it something elese?

OpenStudy (anonymous):

@myininay: Could be. Are you sure in the answer? So So you got A= -(1/16) and B = 1/4. 1/4 is what I got here too but 1/16 the missing coefficient I couldn't solve.

myininaya (myininaya):

i'm checking it it takes alittle time for me to check sorry

OpenStudy (anonymous):

That's fine. Take your time.

OpenStudy (phi):

for undetermined coefficients, you hypothesize a particular solution that is a linear combination of the RHS plus all its derivatives. For example, 2xsinx --> Axsinx+Bcosx (if I did that right) xe^3x --> Cxe^3x + De^3x plug all of that into the differential eq, and match coeff to the RHS

OpenStudy (anonymous):

2xsinx --> Axsinx+Bcosx I was wondering whether it had to be xCosx as well.

OpenStudy (anonymous):

@phi I know. That's why I asked people to solve only with 2xsinx on the right. What I want to know is how to handle coefficient B. I know that for general particular slution I need to add both particular soultions.

myininaya (myininaya):

yes thats right i got it right :)

OpenStudy (anonymous):

Hold on! The way I laearned is that for a 2XsinX the particular solution is: (AX+B)cosX + (AX+B)sinX.

OpenStudy (anonymous):

myininaya Your particular solution for 2XsinX is like mine: (AX+B)cosX + (AX+B)sinX?

OpenStudy (anonymous):

The way I learned was if A CosX + B sin X wasn't going to work, then multiply it by X (possibly more than once9.

myininaya (myininaya):

:)

OpenStudy (anonymous):

@estudier That is if term in particular guess is similar to the homogeneous solution.

myininaya (myininaya):

check those pdfs aarnes

myininaya (myininaya):

the first one shows how i got the particular solution and the 2nd pdf shows me checking

myininaya (myininaya):

it was tedious every move matters that is why it took me so long

OpenStudy (phi):

OK, here's just yp= Axsinx + Bcosx

OpenStudy (phi):

y'= Axcosx + Asinx - Bsinx

OpenStudy (anonymous):

@ I'm checking them right now myininaya.

OpenStudy (phi):

y''= Acosx - Axsinx + Acosx-Bcosx

OpenStudy (phi):

plug those into y'' + 9y= 2xsinx you should get A= 1/4 and B= -1/16

myininaya (myininaya):

and phi is probably right you don't need that whole general particular solution we used but as long as you do things correctly it doesn't matter as long as we have the parts that do matter

OpenStudy (anonymous):

I guess if u do these all the time u learn where the shortcuts are....

myininaya (myininaya):

yay phi! :)

myininaya (myininaya):

we got the same thing but i took a longer route

OpenStudy (anonymous):

Educational.....

OpenStudy (phi):

thanks...

OpenStudy (anonymous):

So is op happy?

myininaya (myininaya):

op?

OpenStudy (anonymous):

original poster

myininaya (myininaya):

oh stayed tuned...

OpenStudy (anonymous):

Whit! @myinimaya: Is this solution to homgeneous eq: http://www.mediafire.com/i/?4ek64ieay6t3l2m

myininaya (myininaya):

whit?

OpenStudy (anonymous):

Wait. :)

OpenStudy (phi):

Yep. That's what wolfram says...

myininaya (myininaya):

looks good

OpenStudy (phi):

Here's how I would do the exponential part: the derivatives of y= \( x e^{3x} \) are \[ y'=3xe^{3x} + e^{3x}\] \[y''=3xe^{3x} + 3e^{3x}+3e^{3x}\] if we keep going, we will see that all the derivatives will be, neglecting constant coefficients, of the form \[ C xe^{3x} + De^{3x} \] C and D to be determined. Assume \( yp=C xe^{3x} + De^{3x} \) plug yp into the differential equation, group like terms, and solve for C and D. If you want I'll grind through it...

OpenStudy (anonymous):

Maybe i's late but here's what a I got: http://www.mediafire.com/i/?juqxn6ro5979oex I'm guessing it's wrong since my only coefficients for first particular solution are A and B. I didn't have C and D. The reason I didn't had them is because of this website: http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx#Second_UnderCoeff_Ex6b Your thoughts?

OpenStudy (phi):

Your yp2 solution looks good. Your starting yp1 does not look right. Each term must have its own coefficient. Instead of (Ax + B)(A cos(x) + B sin(x)) you should use yp1= A x cos(x) + B x sin(x) + C cos(x) + D sin(x)

OpenStudy (phi):

To find an assumed solution, find all the derivates of the RHS= 2 x sin(x) We can ignore the constant 2 (we will be using undetermined coeffs) x sin(x) x cos(x) + sin(x) -x sin(x) + cos(x) + cos(x) it should be clear that if we continue taking derivatives, we will have only terms sin(x), cos(x), x sin(x), x cos(x) neglecting any constant coefficients. So assume yp1= A x cos(x) + B x sin(x) + C cos(x) + D sin(x) Let's see how it turns out. Take the derivative of our assumed solution yp1: yp1= A x cos(x) + B x sin(x) + C cos(x) + D sin(x) yp1' = -A x sin(x) + A cos(x) +B x cos(x) + B sin(x) -C sin(x) + D cos(x) yp1''= -A x cos(x) - A sin(x) - A sin(x) - B x sin(x) + B cos(x) + B cos(x) - C cos(x) - D sin(x) y'' +9y = 2x sin(x) (-A+9A) x cos(x) +(-B+9B) x sin(x) + (2B -C+9C) cos(x) + (-2A-D+9D) sin(x)= 2x sin(x) 8A=0, A=0 8B= 2, B= 1/4 2B+8C=0, C= -1/16 8D=0, D=0 so yp= (1/4) x sin(x) - (1/16) cos(x)

OpenStudy (phi):

As a check, we can use wolfram http://www.wolframalpha.com/input/?i=solve+y%27%27+%2B+9y+%3D+2xsinx

OpenStudy (anonymous):

So assume yp1= A x cos(x) + B x sin(x) + C cos(x) + D sin(x) The key part, same as op but with constants absorbed.

OpenStudy (phi):

aarnes started with (Ax +B)(A cos(x) + Bsin(x)) see http://www.mediafire.com/i/?juqxn6ro5979oex

OpenStudy (anonymous):

yp1= A x cos(x) + B x sin(x) + C cos(x) + D sin(x) -> cosx(Ax +C) + Sinx(Bx+D) Difference is absorbed constants.

OpenStudy (anonymous):

So if there is a polynomial times a trig function we use particular solution AX*cosX + BX*sinX + CcosX + D sinX? Just how did we get to here, phi?

OpenStudy (phi):

I tried to explain it in the above post. Another way, is starting with your (Ax +B)(A cos(x) + Bsin(x)) multiply it out to get A^2 x cos(x) +ABcos(x) +ABx sin(x) +B^2sin(x) and then rename the coeffs. They are all different.

OpenStudy (anonymous):

rename the coeffs. They are all different. Yes, I think that was the cause of the problem (confusing).

OpenStudy (anonymous):

Oh I get it now. From (AX+B)(CcosX+DsinX) we get ACXcosX +ADXsinX+BCcosX+BDsinX then : AXcosX+BXsinX + CcosX + DsinX

OpenStudy (phi):

Yes. That works.

OpenStudy (anonymous):

Great. then what does this website talks about: http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx#Second_UnderCoeff_Ex6b

OpenStudy (phi):

Quoting "So, to avoid this we will do the same thing that we did in the previous example. Everywhere we see a product of constants we will rename it and call it a single constant." He is doing exactly what you just did. Rename the coeffs.

OpenStudy (anonymous):

Yes, I think so too, he is usually pretty reliable, that fellow...

OpenStudy (anonymous):

Aaah. Yes he did. My mistake then. I was fast reading and didn't see that he introduced D E and F constants. Oh man. :D

OpenStudy (anonymous):

Well, we can call this topic closed! :) Thank you all for help.